The time period of satellite of earth is 7 hours.if separation b/w the earth and satellite is increased to twice the previous value. What will be it`s new time period?????

Grade:11

1 Answers

Arun
6 years ago
According to Kepler’s law of planetary motion T2 =K. R3  T2= T1 (R2/R)3/2 = 7 x [2R/R]3/2= 7 x 2\sqrt2 = 14\sqrt2 hours